Circular No. 7355

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999em IN NGC 1637 L. Wang, University of Texas; D. Baade, European Southern Obser-

vatory (ESO); F. Ma, P. Hoeflich, and J. C. Wheeler, University of Texas; and C. Fransson, Stockholm Observatory, report: “Spectropolarimetry of SN 1999em was obtained using the ESO Very Large Telescope (VLT) on 1999 Nov. 3 and 2000 Jan. 9 UT (ranges 360–880 and 416–950 nm for the Nov. 3 and Jan. 9 data, respectively). The Nov. 3 data show weak continuum polarization of ∼ 0.25 percent and hints of depolarization across the Hα P-Cyg profile. The Jan. 9 data show clearly higher continuum polarization to ∼ 0.5 percent and strong modulations across most of the strong absorption/emission features. Notably, nearly complete depolarizations are observed at the emission peaks of Hα and He/Na I 587.6-nm lines. Spectropolarimetry (range 415–780 nm) was also obtained using the 2.1-m telescope of the McDonald Observatory durng Jan. 7–10; the results are in agreement with the VLT data taken on Jan. 9. The Nov. 3 VLT data are in qualitative agreement with previous reports (IAUC 7305). The time variation of the polarization and the change across emission lines suggest that the polarization is due to scattering by aspherical supernova ejecta and not interstellar dust. The increasing polarization at later times implies higher asphericity deeper into the core of the supernova. Continuous monitoring at later epochs will provide valuable information on the explosion mechanisms of type-II supernovae.”

SAX J2103.5+4545 A. Baykal, Middle East Technical University; M. Stark, Marietta Col-

lege; and J. H. Swank, Goddard Space Flight Center, report: “Timing analysis of RXTE PCA observations of the transient x-ray pulsar SAX J2103.5+4545 (Hulleman et al. 1998, A.Ap. 337, L25), in outburst since 1999 Oct. 25, shows an eccentric orbit with P = 12.68 ± 0.25 days, e = 0.4 ± 0.2, and projected a = 72 ± 6 s. The pulse period of 358.62171(88) is consistent with the 1997 period, but the spin-up rate during 1999 Nov. 19–27 was 2.50(15) ∼ 10^{−13} Hz s^{−1}, at a flux of 15–20 mCrab, implying (for a 10^{12} Gauss field) a luminosity of 8 × 10^{35} erg s^{−1} at a distance of 4 kpc — probably beyond the star HD 200709, which was found just outside the BeppoSAX error circle. The RXTE ASM record shows that the x-ray flux is sometimes modulated on a time scale of 13 days, and as much as 20 percent during the 1997 outburst. Deeper searches for an optical counterpart are urged.”

2000 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green